# Ringdown in Vaidya spacetimes: time-dependent frequencies, Penrose limit and time-domain analyses

### v2

@@ III. Penrose limit and Ringdown analysis in a Vaidya spacetime with a constant accretion rate
$\mathrm{d}s^{2}=-\left(1-\frac{2M(\mathcal{V})}{r}\right)\mathrm{d}\mathcal{V}^{2}+2\mathrm{d}\mathcal{V}\mathrm{d}r+r^{2}\left(\mathrm{d}\theta^{2}+\sin^{2}\theta\mathrm{d}\phi^{2}\right).$ (46)
@@
$M(\mathcal{V})=M_{1}+M^{\prime}\mathcal{V}$ (47)
with $M^{\prime}={\rm const.}>0$. The frequency domain analyses for the static conformal metric have been done in Ref. [34]. We reanalyze it in the time domain and compare the results with the frequency domain analyses.
@@
$x$ $=$ $\frac{r}{2M(\mathcal{V})},$ (48)
$T$ $=$ $\int\frac{\mathrm{d}\mathcal{V}}{2M(\mathcal{V})}-x_{*}(x)$ (49)
$=$ $\frac{1}{2M^{\prime}}\ln\left|1+M^{\prime}\frac{\mathcal{V}}{M_{1}}\right|-x_{*}(x),$ (50)
@@
$x_{*}(x)=\int\frac{\mathrm{d}x}{\tilde{f}(x)}$ (51)
@@
$\tilde{f}(x)=1-\frac{1}{x}-4M^{\prime}x.$ (52)
@@
$\mathrm{d}s^{2}=g^{(0)}_{\mu\nu}\mathrm{d}\tilde{x}^{\mu}\mathrm{d}\tilde{x}^{\nu}:=4M^{2}(\mathcal{V})\left[-\tilde{f}(x)\mathrm{d}T^{2}+\frac{1}{\tilde{f}(x)}\mathrm{d}x^{2}+x^{2}\mathrm{d}\Omega^{2}\right],$ (53)
@@
$\tilde{g}^{(0)}_{\mu\nu}\mathrm{d}\tilde{x}^{\mu}\mathrm{d}\tilde{x}^{\nu}=-\tilde{f}(x)\mathrm{d}T^{2}+\frac{1}{\tilde{f}(x)}\mathrm{d}x^{2}+x^{2}\mathrm{d}\Omega^{2}.$ (54)
@@
$x=x_{\mp}:=\frac{1\mp\sqrt{1-16M^{\prime}}}{8M^{\prime}}.$ (55)
@@ A. Penrose limit analysis
$x_{\rm c}=\frac{1-\sqrt{1-12M^{\prime}}}{4M^{\prime}},$ (56)
where we took the smaller root associated with the event horizon. Then the quasinormal mode frequency of the fundamental mode $n=0$ can be estimated as
$\Omega^{\rm Re}_{\rm PL}$ $=$ $\frac{\sqrt{\tilde{f}(x_{\rm c})}}{x_{c}}\left(\ell+\frac{1}{2}\right),$ (57)
$\Omega^{\rm Im}_{\rm PL}$ $=$ $-\frac{1}{2}\sqrt{\frac{\tilde{f}(x_{\rm c})^{2}}{x_{\rm c}^{2}}-\frac{1}{2}\tilde{f}(x_{\rm c})\tilde{f}^{\prime\prime}(x_{\rm c})}.$ (58)
@@
$\Omega^{\rm Re}_{\rm PL}$ $=$ $-\frac{2\sqrt{6\sqrt{1-12M^{\prime}}-3}}{3\left(\sqrt{1-12M^{\prime}}-1\right)}M^{\prime}(2\ell+1),$ (59)
$\Omega^{\rm Im}_{\rm PL}$ $=$ $-\frac{\sqrt{6M^{\prime}\left(24M^{\prime}-3\sqrt{1-12M^{\prime}}-4\right)+\sqrt{1-12M^{\prime}}+1}}{3\sqrt{6}}.$ (60)
@@
$\mathcal{R}_{\rm PL}:=\left|\frac{\Omega^{\rm Im}_{\rm PL}}{\Omega^{\rm Re}_{\rm PL}}\right|=\frac{\left(\sqrt{1-12M^{\prime}}-1\right)\sqrt{6M^{\prime}\left(24M^{\prime}-3\sqrt{1-12M^{\prime}}-4\right)+\sqrt{1-12M^{\prime}}+1}}{6(2\ell+1)M^{\prime}\sqrt{4\sqrt{1-12M^{\prime}}-2}}.$ (61)
The value of $1-(2\ell+1)\mathcal{R}_{\rm PL}$ is shown as a function of $M^{\prime}$ in Fig. 1. The ratio $(2\ell+1)\mathcal{R}_{\rm PL}$ decreases with the value of $M^{\prime}$ compared to 1 for the Schwarzschild case.
Figure 1: The value of $1-(2\ell+1)\mathcal{R}_{\rm PL}$ is shown as a function of $M^{\prime}$.
@@ B. Frequency domain analyses
$\nabla_{\mu}\nabla^{\mu}\Phi=\left(\tilde{\nabla}_{\mu}\tilde{\nabla}^{\mu}-\frac{\tilde{R}}{6}\right)\tilde{\Phi}=0,$ (63)
@@
$\tilde{\Phi}={\rm e}^{-\mathrm{i}\tilde{\Omega}T}\frac{\psi^{(0)}(x)}{x}Y_{\ell m}(\theta,\phi),$ (64)
@@
$\left[\frac{\mathrm{d}^{2}}{\mathrm{d}x_{*}^{2}}+\tilde{\Omega}^{2}-\tilde{f}(x)\left(\frac{\ell(\ell+1)}{x^{2}}+\frac{1}{x^{3}}\right)\right]\psi^{(0)}(x)=0.$ (65)
@@
$A_{\mu}\mathrm{d}\tilde{x}^{\mu}={\rm e}^{-\mathrm{i}\tilde{\Omega}T}\psi^{(1)}(x)\epsilon_{i}^{~j}\partial_{j}Y_{\ell m}\mathrm{d}\tilde{x}^{i},$ (66)
@@
$\left[\frac{\mathrm{d}^{2}}{\mathrm{d}x_{*}^{2}}+\tilde{\Omega}^{2}-\tilde{f}(x)\left(\frac{\ell(\ell+1)}{x^{2}}\right)\right]\psi^{(1)}(x)=0.$ (67)
@@
$\tilde{h}_{\mu\nu}\mathrm{d}\tilde{x}^{\mu}\mathrm{d}\tilde{x}^{\nu}$ $=$ ${\rm e}^{-\mathrm{i}\tilde{\Omega}T}\Biggl[2h_{0}(x)\left(-\frac{1}{\sin\theta}\partial_{\phi}Y_{\ell m}\mathrm{d}T\mathrm{d}\theta+\sin\theta\partial_{\theta}Y_{\ell m}\mathrm{d}T\mathrm{d}\phi\right)$ (69)
@@
$\tilde{R}_{\mu\nu}-2\tilde{\nabla}_{\mu}\tilde{\nabla}_{\nu}\ln M-\tilde{g}_{\mu\nu}\tilde{\nabla}_{\alpha}\tilde{\nabla}^{\alpha}\ln M+2\tilde{\nabla}_{\mu}\ln M\tilde{\nabla}_{\nu}\ln M-2\tilde{g}_{\mu\nu}\tilde{\nabla}_{\alpha}\ln M\tilde{\nabla}^{\alpha}\ln M=8\pi T_{\mu\nu}$ (70)
@@
$T_{\mu\nu}\mathrm{d}\tilde{x}^{\mu}\mathrm{d}\tilde{x}^{\nu}=\frac{M^{\prime}}{4\pi x^{2}}\left[\mathrm{d}T^{2}+\frac{2}{\tilde{f}(x)}\mathrm{d}T\mathrm{d}x+\frac{\mathrm{d}x^{2}}{\tilde{f}(x)^{2}}\right].$ (71)
@@
$F(x)=-\frac{x^{5/4}}{f(x)^{3/4}}\exp\left(\frac{{\rm ArcTan}\left(\frac{8M^{\prime}x-1}{\sqrt{16M^{\prime}-1}}\right)}{2\sqrt{16M^{\prime}-1}}\right),$ (72)
@@
$\left[\frac{\mathrm{d}^{2}}{\mathrm{d}x_{*}^{2}}+\left(\tilde{\Omega}+2\mathrm{i}M^{\prime}\right)^{2}-\tilde{f}(x)\left(\frac{\ell(\ell+1)}{x^{2}}-\frac{3}{x^{3}}\right)\right]\psi^{(2)}(x)=0.$ (73)
@@
$\left[\frac{\mathrm{d}^{2}}{\mathrm{d}x_{*}^{2}}+\hat{\Omega}^{2}-\tilde{f}(x)\left(\frac{\ell(\ell+1)}{x^{2}}-\frac{s^{2}-1}{x^{3}}\right)\right]\psi^{(s)}(x)=0,$ (74)
where $\hat{\Omega}=\tilde{\Omega}$ for $s=0$ and $s=1$, and $\hat{\Omega}=\tilde{\Omega}+2\mathrm{i}M^{\prime}$ for $s=2$. The master equations can be solved with purely ingoing/outgoing boundary conditions at $x=x_{\mp}$. The results with $M^{\prime}=\pi/2000$ and $\ell=4$ are summarized in Table 1.
Table 1: The real and imaginary quasinormal mode frequencies in the conformal static spacetime of the metric $\tilde{g}_{\mu\nu}$ with $M^{\prime}=\pi/2000$ and $\ell=4$.
@@
To compare the analysis of the Penrose limit geometry with the ringdown waveform in Vaidya spacetimes, we numerically solve the tensor mode master equation in the time domain by using the double null formalism following Refs. [39, 40]. First, let us consider the relation between the double null coordinates $(\mathcal{U},\mathcal{V}$) and the coordinates $(\mathcal{V},r)$ in Eq. (46). In the double null coordinates, the metric form is given as
$\mathrm{d}s^{2}=-{\rm e}^{\sigma(\mathcal{U},\mathcal{V})}\mathrm{d}\mathcal{U}\mathrm{d}\mathcal{V}+r^{2}(\mathcal{U},\mathcal{V})(\mathrm{d}\theta^{2}+\sin^{2}\theta\mathrm{d}\phi^{2}).$ (75)
@@
$\mathrm{d}\mathcal{U}=A(\mathcal{V},r)\mathrm{d}\mathcal{V}-B(\mathcal{V},r)\mathrm{d}r,$ (76)
@@
$A=\frac{1}{2}\left(1-\frac{2M(\mathcal{V})}{r}\right)B.$ (77)
@@
$\mathrm{d}r=-\frac{1}{B}\mathrm{d}\mathcal{U}+\frac{1}{2}\left(1-\frac{2M(\mathcal{V})}{r}\right)\mathrm{d}\mathcal{V},$ (78)
@@
$\frac{\partial r}{\partial\mathcal{U}}$ $=$ $-\frac{1}{B}=-\frac{1}{2}{\rm e}^{\sigma(\mathcal{U},\mathcal{V})},$ (79)
$\frac{\partial r}{\partial\mathcal{V}}$ $=$ $\frac{1}{2}\left(1-\frac{2M(\mathcal{V})}{r}\right),$ (80)
$\frac{\partial^{2}r}{\partial\mathcal{U}\partial\mathcal{V}}$ $=$ $\frac{\partial}{\partial\mathcal{U}}\left[\frac{1}{2}\left(1-\frac{2M(\mathcal{V})}{r}\right)\right].$ (81)
We numerically solve Eq. (81) with the boundary condition Eq. (80) on the outermost grid points along $\mathcal{U}={\rm const.}$ line.
@@
$\Phi=\frac{\Psi^{(0)}(\mathcal{U},\mathcal{V})}{r(\mathcal{U},\mathcal{V})}Y_{\ell m}(\theta,\phi).$ (82)
@@
$\frac{\partial^{2}\Psi^{(0)}}{\partial\mathcal{U}\partial\mathcal{V}}=-\frac{1}{2r^{2}}\frac{\partial r}{\partial\mathcal{U}}\left[\ell(\ell+1)+\frac{2M(\mathcal{V})}{r}\right]\Psi^{(0)}.$ (83)
@@
$A_{\mu}\mathrm{d}x^{\mu}=\Psi^{(1)}(\mathcal{U},\mathcal{V})\epsilon_{i}^{~j}\partial_{j}Y_{\ell m}\mathrm{d}x^{i},$ (84)
@@
$\frac{\partial^{2}\Psi^{(1)}}{\partial\mathcal{U}\partial\mathcal{V}}=-\frac{1}{2r^{2}}\frac{\partial r}{\partial\mathcal{U}}\left[\ell(\ell+1)\right]\Psi^{(1)}.$ (85)
@@
$h_{\mu\nu}\mathrm{d}x^{\mu}\mathrm{d}x^{\nu}$ $=$ $2H_{0}(\mathcal{U},\mathcal{V})\left(-\frac{1}{\sin\theta}\partial_{\phi}Y_{\ell m}\mathrm{d}\mathcal{U}\mathrm{d}\theta+\sin\theta\partial_{\theta}Y_{\ell m}\mathrm{d}\mathcal{U}\mathrm{d}\phi\right)$ (87)
@@
$\Psi^{(2)}(\mathcal{U},\mathcal{V})=-\frac{4{\rm e}^{-\sigma}}{(\ell-1)(\ell+2)M_{1}}\left[2\left(\frac{\partial r}{\partial\mathcal{V}}H_{0}-\frac{\partial r}{\partial\mathcal{U}}H_{1}\right)+r\left(\frac{\partial H_{1}}{\partial\mathcal{U}}-\frac{\partial H_{0}}{\partial\mathcal{V}}\right)\right],$ (88)
@@
$\frac{\partial^{2}\Psi^{(2)}}{\partial\mathcal{U}\partial\mathcal{V}}=-\frac{1}{2r^{2}}\frac{\partial r}{\partial\mathcal{U}}\left[\ell(\ell+1)-\frac{6M(\mathcal{V})}{r}\right]\Psi^{(2)}.$ (89)
@@
$H_{0}$ $=$ $\frac{M_{1}}{2}\left(\frac{\partial r}{\partial\mathcal{U}}\Psi^{(2)}+r\frac{\partial\Psi^{(2)}}{\partial\mathcal{U}}\right),$ (90)
$H_{1}$ $=$ $\frac{M_{1}}{2}\left(\frac{\partial r}{\partial\mathcal{V}}\Psi^{(2)}+r\frac{\partial\Psi^{(2)}}{\partial\mathcal{V}}\right).$ (91)
@@
$\frac{\partial^{2}\Psi^{(s)}}{\partial\mathcal{U}\partial\mathcal{V}}=-\frac{1}{2r^{2}}\frac{\partial r}{\partial\mathcal{U}}\left[\ell(\ell+1)-(s^{2}-1)\frac{2M(\mathcal{V})}{r}\right]\Psi^{(s)}$ (92)
@@
For the Schwarzschild spacetime, with the mass $M$, the master equation (92) can be reduced to the form:
$\left[\frac{\mathrm{d}^{2}}{\mathrm{d}r_{*}^{2}}+\omega_{\rm Sch}^{2}-\left(1-\frac{2M}{r}\right)\left(\frac{\ell(\ell+1)}{r^{2}}-\frac{s^{2}-1}{r^{3}}\right)\right]\psi_{\rm Sch}^{(s)}(x)=0,$ (93)
@@
where $r_{*}$ is the tortoise coordinate, and we have introduced the frequency $\omega_{\rm Sch}$. For later convenience, we show the values of $\omega^{\rm Im/Re}_{\rm PL}$ and numerically calculated values of $\omega_{\rm Sch}^{\rm Im/Re}$ of the fundamental mode in the Schwarzschild spacetime with $\ell=4$ in Table 2. We also calculate the numerical factor $\mathcal{C}:=(2\ell+1)\omega^{\rm Im}/\omega^{\rm Re}$, which gives $1$ for the Penrose-limit analysis. The values of $\mathcal{C}$ can be regarded as correction factors for the value of $(2\ell+1)\omega^{\rm Im}/\omega^{\rm Re}$ of each field needed for comparing the Penrose limit analysis with the numerical ringdown analyses.

### v3

@@ III. Penrose limit and Ringdown analysis in a Vaidya spacetime with a constant accretion rate
$\mathrm{d}s^{2}=-\left(1-\frac{2M(\mathcal{V})}{r}\right)\mathrm{d}\mathcal{V}^{2}+2\mathrm{d}\mathcal{V}\mathrm{d}r+r^{2}\left(\mathrm{d}\theta^{2}+\sin^{2}\theta\mathrm{d}\phi^{2}\right).$ (3.1)
@@
$M(\mathcal{V})=M_{1}+M^{\prime}\mathcal{V}$ (3.2)
with $M^{\prime}={\rm const.}>0$. The frequency-domain analyses for the static conformal metric have been done in Ref. [34]. We reanalyze it in the time domain and compare the results with the frequency domain analyses.
@@
$x$ $=$ $\frac{r}{2M(\mathcal{V})},$ (3.3)
$T$ $=$ $\int\frac{\mathrm{d}\mathcal{V}}{2M(\mathcal{V})}-x_{*}(x)$ (3.4)
$=$ $\frac{1}{2M^{\prime}}\ln\left|1+M^{\prime}\frac{\mathcal{V}}{M_{1}}\right|-x_{*}(x),$ (3.5)
@@
$x_{*}(x)=\int\frac{\mathrm{d}x}{\tilde{f}(x)}$ (3.6)
@@
$\tilde{f}(x)=1-\frac{1}{x}-4M^{\prime}x.$ (3.7)
@@
$\mathrm{d}s^{2}=g^{(0)}_{\mu\nu}\mathrm{d}\tilde{x}^{\mu}\mathrm{d}\tilde{x}^{\nu}:=4M^{2}(\mathcal{V})\left[-\tilde{f}(x)\mathrm{d}T^{2}+\frac{1}{\tilde{f}(x)}\mathrm{d}x^{2}+x^{2}\mathrm{d}\Omega^{2}\right],$ (3.8)
@@
$\tilde{g}^{(0)}_{\mu\nu}\mathrm{d}\tilde{x}^{\mu}\mathrm{d}\tilde{x}^{\nu}=-\tilde{f}(x)\mathrm{d}T^{2}+\frac{1}{\tilde{f}(x)}\mathrm{d}x^{2}+x^{2}\mathrm{d}\Omega^{2}.$ (3.9)
@@
$x=x_{\mp}:=\frac{1\mp\sqrt{1-16M^{\prime}}}{8M^{\prime}}.$ (3.10)
@@ A. Penrose limit analysis
$x_{\rm c}=\frac{1-\sqrt{1-12M^{\prime}}}{4M^{\prime}},$ (3.11)
where we took the smaller root associated with the event horizon. Then the QNM frequency of the fundamental mode $n=0$ can be estimated as
$\Omega^{\rm Re}_{\rm PL}$ $=$ $\frac{\sqrt{\tilde{f}(x_{\rm c})}}{x_{c}}\left(\ell+\frac{1}{2}\right),$ (3.12)
$\Omega^{\rm Im}_{\rm PL}$ $=$ $-\frac{1}{2}\sqrt{\frac{\tilde{f}(x_{\rm c})^{2}}{x_{\rm c}^{2}}-\frac{1}{2}\tilde{f}(x_{\rm c})\tilde{f}^{\prime\prime}(x_{\rm c})}.$ (3.13)
@@
$\Omega^{\rm Re}_{\rm PL}$ $=$ $-\frac{2\sqrt{6\sqrt{1-12M^{\prime}}-3}}{3\left(\sqrt{1-12M^{\prime}}-1\right)}M^{\prime}(2\ell+1),$ (3.14)
$\Omega^{\rm Im}_{\rm PL}$ $=$ $-\frac{\sqrt{6M^{\prime}\left(24M^{\prime}-3\sqrt{1-12M^{\prime}}-4\right)+\sqrt{1-12M^{\prime}}+1}}{3\sqrt{6}}.$ (3.15)
@@
$\mathcal{R}_{\rm PL}:=\left|\frac{\Omega^{\rm Im}_{\rm PL}}{\Omega^{\rm Re}_{\rm PL}}\right|=\frac{\left(\sqrt{1-12M^{\prime}}-1\right)\sqrt{6M^{\prime}\left(24M^{\prime}-3\sqrt{1-12M^{\prime}}-4\right)+\sqrt{1-12M^{\prime}}+1}}{6(2\ell+1)M^{\prime}\sqrt{4\sqrt{1-12M^{\prime}}-2}}.$ (3.16)
The value of $1-(2\ell+1)\mathcal{R}_{\rm PL}$ is shown as a function of $M^{\prime}$ in Fig. 1. In Fig. 1, the specific values of $M^{\prime}$ indicated by the vertical lines correspond to the values of $M^{\prime}$ (Sec. III) and $M^{\prime}_{\rm max}$ (Sec. IV) for specific examples shown in the subsequent sections. The ratio $(2\ell+1)\mathcal{R}_{\rm PL}$ decreases with the value of $M^{\prime}$ compared to 1 for the Schwarzschild case.
Figure 1: The value of $1-(2\ell+1)\mathcal{R}_{\rm PL}$ is shown as a function of $M^{\prime}$. The specific values of $M^{\prime}$ indicated by the vertical lines correspond to the values of $M^{\prime}$ (Sec. III) and $M^{\prime}_{\rm max}$ (Sec. IV) for specific examples shown in the subsequent sections.
@@ B. Frequency domain analyses
$\nabla_{\mu}\nabla^{\mu}\Phi=\left(\tilde{\nabla}_{\mu}\tilde{\nabla}^{\mu}-\frac{\tilde{R}}{6}\right)\tilde{\Phi}=0,$ (3.18)
@@
$\tilde{\Phi}={\rm e}^{-\mathrm{i}\tilde{\Omega}T}\frac{\psi^{(0)}(x)}{x}Y_{\ell m}(\theta,\phi),$ (3.19)
@@
$\left[\frac{\mathrm{d}^{2}}{\mathrm{d}x_{*}^{2}}+\tilde{\Omega}^{2}-\tilde{f}(x)\left(\frac{\ell(\ell+1)}{x^{2}}+\frac{1}{x^{3}}\right)\right]\psi^{(0)}(x)=0.$ (3.20)
@@
$A_{\mu}\mathrm{d}\tilde{x}^{\mu}={\rm e}^{-\mathrm{i}\tilde{\Omega}T}\psi^{(1)}(x)\epsilon_{i}^{~j}\partial_{j}Y_{\ell m}\mathrm{d}\tilde{x}^{i},$ (3.21)
@@
$\left[\frac{\mathrm{d}^{2}}{\mathrm{d}x_{*}^{2}}+\tilde{\Omega}^{2}-\tilde{f}(x)\left(\frac{\ell(\ell+1)}{x^{2}}\right)\right]\psi^{(1)}(x)=0.$ (3.22)
@@
$\tilde{h}_{\mu\nu}\mathrm{d}\tilde{x}^{\mu}\mathrm{d}\tilde{x}^{\nu}$ $=$ ${\rm e}^{-\mathrm{i}\tilde{\Omega}T}\Biggl[2h_{0}(x)\left(-\frac{1}{\sin\theta}\partial_{\phi}Y_{\ell m}\mathrm{d}T\mathrm{d}\theta+\sin\theta\partial_{\theta}Y_{\ell m}\mathrm{d}T\mathrm{d}\phi\right)$ (3.24)
@@
$\tilde{R}_{\mu\nu}-2\tilde{\nabla}_{\mu}\tilde{\nabla}_{\nu}\ln M-\tilde{g}_{\mu\nu}\tilde{\nabla}_{\alpha}\tilde{\nabla}^{\alpha}\ln M+2\tilde{\nabla}_{\mu}\ln M\tilde{\nabla}_{\nu}\ln M-2\tilde{g}_{\mu\nu}\tilde{\nabla}_{\alpha}\ln M\tilde{\nabla}^{\alpha}\ln M=8\pi T_{\mu\nu}$ (3.25)
@@
$T_{\mu\nu}\mathrm{d}\tilde{x}^{\mu}\mathrm{d}\tilde{x}^{\nu}=\frac{M^{\prime}}{4\pi x^{2}}\left[\mathrm{d}T^{2}+\frac{2}{\tilde{f}(x)}\mathrm{d}T\mathrm{d}x+\frac{\mathrm{d}x^{2}}{\tilde{f}(x)^{2}}\right].$ (3.26)
@@
$F(x)=-\frac{x^{5/4}}{f(x)^{3/4}}\exp\left(\frac{{\rm arctan}\left(\frac{8M^{\prime}x-1}{\sqrt{16M^{\prime}-1}}\right)}{2\sqrt{16M^{\prime}-1}}\right),$ (3.27)
@@
$\left[\frac{\mathrm{d}^{2}}{\mathrm{d}x_{*}^{2}}+\left(\tilde{\Omega}+2\mathrm{i}M^{\prime}\right)^{2}-\tilde{f}(x)\left(\frac{\ell(\ell+1)}{x^{2}}-\frac{3}{x^{3}}\right)\right]\psi^{(2)}(x)=0.$ (3.28)
@@
$\left[\frac{\mathrm{d}^{2}}{\mathrm{d}x_{*}^{2}}+\hat{\Omega}^{2}-\tilde{f}(x)\left(\frac{\ell(\ell+1)}{x^{2}}-\frac{s^{2}-1}{x^{3}}\right)\right]\psi^{(s)}(x)=0,$ (3.29)
where $\hat{\Omega}=\tilde{\Omega}$ for $s=0$ and $s=1$, and $\hat{\Omega}=\tilde{\Omega}+2\mathrm{i}M^{\prime}$ for $s=2$. The master equations can be solved with purely ingoing/outgoing boundary conditions at $x=x_{\mp}$. The results obtained by the Leaver’s method [39] (see, e.g., a review [40]) with $M^{\prime}=\pi/2000$ and $\ell=4$ are summarized in Table 1.
Table 1: The real and imaginary QNM frequencies in the conformal static spacetime of the metric $\tilde{g}_{\mu\nu}$ with $M^{\prime}=\pi/2000$ and $\ell=4$.
@@ C. Time-domain analyses
To compare the analysis of the PL geometry with the ringdown waveform in Vaidya spacetimes, we numerically solve the tensor mode master equation in the time domain by using the double null formalism following Refs. [41, 42]. First, let us consider the relation between the double null coordinates $(\mathcal{U},\mathcal{V}$) and the coordinates $(\mathcal{V},r)$ in Eq. (3.1). In the double null coordinates, the metric form is given as
$\mathrm{d}s^{2}=-{\rm e}^{\sigma(\mathcal{U},\mathcal{V})}\mathrm{d}\mathcal{U}\mathrm{d}\mathcal{V}+r^{2}(\mathcal{U},\mathcal{V})(\mathrm{d}\theta^{2}+\sin^{2}\theta\mathrm{d}\phi^{2}).$ (3.30)
@@
$\mathrm{d}\mathcal{U}=A(\mathcal{V},r)\mathrm{d}\mathcal{V}-B(\mathcal{V},r)\mathrm{d}r,$ (3.31)
@@
$A=\frac{1}{2}\left(1-\frac{2M(\mathcal{V})}{r}\right)B.$ (3.32)
@@
$\mathrm{d}r=-\frac{1}{B}\mathrm{d}\mathcal{U}+\frac{1}{2}\left(1-\frac{2M(\mathcal{V})}{r}\right)\mathrm{d}\mathcal{V},$ (3.33)
@@
$\frac{\partial r}{\partial\mathcal{U}}$ $=$ $-\frac{1}{B}=-\frac{1}{2}{\rm e}^{\sigma(\mathcal{U},\mathcal{V})},$ (3.34)
$\frac{\partial r}{\partial\mathcal{V}}$ $=$ $\frac{1}{2}\left(1-\frac{2M(\mathcal{V})}{r}\right),$ (3.35)
$\frac{\partial^{2}r}{\partial\mathcal{U}\partial\mathcal{V}}$ $=$ $\frac{\partial}{\partial\mathcal{U}}\left[\frac{1}{2}\left(1-\frac{2M(\mathcal{V})}{r}\right)\right].$ (3.36)
We numerically solve Eq. (3.36) with the boundary condition Eq. (3.35) on the outermost grid points along $\mathcal{U}={\rm const.}$ line.
@@
$\Phi=\frac{\Psi^{(0)}(\mathcal{U},\mathcal{V})}{r(\mathcal{U},\mathcal{V})}Y_{\ell m}(\theta,\phi).$ (3.37)
@@
$\frac{\partial^{2}\Psi^{(0)}}{\partial\mathcal{U}\partial\mathcal{V}}=-\frac{1}{2r^{2}}\frac{\partial r}{\partial\mathcal{U}}\left[\ell(\ell+1)+\frac{2M(\mathcal{V})}{r}\right]\Psi^{(0)}.$ (3.38)
@@
$A_{\mu}\mathrm{d}x^{\mu}=\Psi^{(1)}(\mathcal{U},\mathcal{V})\epsilon_{i}^{~j}\partial_{j}Y_{\ell m}\mathrm{d}x^{i},$ (3.39)
@@
$\frac{\partial^{2}\Psi^{(1)}}{\partial\mathcal{U}\partial\mathcal{V}}=-\frac{1}{2r^{2}}\frac{\partial r}{\partial\mathcal{U}}\left[\ell(\ell+1)\right]\Psi^{(1)}.$ (3.40)
@@
$h_{\mu\nu}\mathrm{d}x^{\mu}\mathrm{d}x^{\nu}$ $=$ $2H_{0}(\mathcal{U},\mathcal{V})\left(-\frac{1}{\sin\theta}\partial_{\phi}Y_{\ell m}\mathrm{d}\mathcal{U}\mathrm{d}\theta+\sin\theta\partial_{\theta}Y_{\ell m}\mathrm{d}\mathcal{U}\mathrm{d}\phi\right)$ (3.42)
@@
$\Psi^{(2)}(\mathcal{U},\mathcal{V})=-\frac{4{\rm e}^{-\sigma}}{(\ell-1)(\ell+2)M_{1}}\left[2\left(\frac{\partial r}{\partial\mathcal{V}}H_{0}-\frac{\partial r}{\partial\mathcal{U}}H_{1}\right)+r\left(\frac{\partial H_{1}}{\partial\mathcal{U}}-\frac{\partial H_{0}}{\partial\mathcal{V}}\right)\right],$ (3.43)
@@
$\frac{\partial^{2}\Psi^{(2)}}{\partial\mathcal{U}\partial\mathcal{V}}=-\frac{1}{2r^{2}}\frac{\partial r}{\partial\mathcal{U}}\left[\ell(\ell+1)-\frac{6M(\mathcal{V})}{r}\right]\Psi^{(2)}.$ (3.44)
@@
$H_{0}$ $=$ $\frac{M_{1}}{2}\left(\frac{\partial r}{\partial\mathcal{U}}\Psi^{(2)}+r\frac{\partial\Psi^{(2)}}{\partial\mathcal{U}}\right),$ (3.45)
$H_{1}$ $=$ $\frac{M_{1}}{2}\left(\frac{\partial r}{\partial\mathcal{V}}\Psi^{(2)}+r\frac{\partial\Psi^{(2)}}{\partial\mathcal{V}}\right).$ (3.46)
@@
$\frac{\partial^{2}\Psi^{(s)}}{\partial\mathcal{U}\partial\mathcal{V}}=-\frac{1}{2r^{2}}\frac{\partial r}{\partial\mathcal{U}}\left[\ell(\ell+1)-(s^{2}-1)\frac{2M(\mathcal{V})}{r}\right]\Psi^{(s)}$ (3.47)
@@
For the Schwarzschild spacetime, with the mass $M$, the master equation (3.47) can be reduced to the form
$\left[\frac{\mathrm{d}^{2}}{\mathrm{d}r_{*}^{2}}+\omega_{\rm Sch}^{2}-\left(1-\frac{2M}{r}\right)\left(\frac{\ell(\ell+1)}{r^{2}}-\frac{s^{2}-1}{r^{3}}\right)\right]\psi_{\rm Sch}^{(s)}(x)=0,$ (3.48)
@@
where $r_{*}$ is the tortoise coordinate, and we have introduced the frequency $\omega_{\rm Sch}$. For later convenience, we show the values of $\omega^{\rm Im/Re}_{\rm PL}$ and numerically calculated values of $\omega_{\rm Sch}^{\rm Im/Re}$ of the fundamental mode in the Schwarzschild spacetime with $\ell=4$ in Table 2. We also calculate the numerical factor $\mathcal{C}:=(2\ell+1)|\omega^{\rm Im}/\omega^{\rm Re}|$, which gives $1$ for the Penrose-limit analysis. The values of $\mathcal{C}$ can be regarded as correction factors for the value of $(2\ell+1)|\omega^{\rm Im}/\omega^{\rm Re}|$ of each field needed for comparing the PL analysis with the numerical ringdown analyses. It should be noted that we focus on the fundamental modes in this paper, and the values of $\mathcal{C}$ are dependent on the overtone number.